The rest mass of an electron is 9.1 x


10^-31 kg. What is the relativistic mass if its velocity is 55 % of the speed


of light?

1
Expert's answer
2022-02-09T06:27:14-0500

Explanation & Calculation


  • The relativistic mass is given by the following formula and upon substitution of the given values into it, you can get the answer.

"\qquad\qquad
\begin{aligned}
\small M_{relativistic}&=\small \frac{m_{rest}}{\sqrt{1-\large\frac{v^2}{c^2}}}
\end{aligned}"

  • Here the speed is

"\qquad\qquad
\begin{aligned}
\small v&=\small \frac{55}{100}c\\
\implies \small \frac{v^2}{c^2}&=\small\frac{55^2}{100^2}
\end{aligned}"

  • Then the required answer is

"\qquad\qquad
\begin{aligned}
\small M_r&=\small \frac{9.1\times10^{-31}\,kg}{\sqrt{1-\frac{55^2}{100^2}}}\\
&=\small 1.1\times10^{-30}\,kg\cdots(\text{increased})
\end{aligned}"